OpenStudy (anonymous):

What is the value of "c" in the quadratic equation 3x^2 + 5x + 7 = 0? 3 5 7

OpenStudy (anonymous):

it is C: 7

OpenStudy (anonymous):

Write the quadratic equation in general form. (x - 3)2 = 0 x^2 + 9 = 0 x^2 + 6x + 9 = 0 x^2 - 6x + 9 = 0

OpenStudy (anonymous):

simply use the formula (a+b)^2=a^2+b^2+2ab or simply (x-3)x (x-3)=x^2-6x+9=0 where a=1,b=-6,c=9